suppose we observed an economy in which changes in the money supply produce no changes whatever in nominal GDP. what could we conclude about velocity
Answer #1

Ans. We know that M × V = Nominal GDP

Now if there is no change in Nominal GDP due to change in money supply , this would mean that Velocity of money would be inversely and proportionally related to money supply.

For eg if initially , 1 × 1 = 1

Then if money supply increased to 5 , then

( 1 × 5 ) × 1 ( 1 ÷ 5 ) = 1

This means that Nominal GDP will not change.

Because Velocity of money is equal to 1 divided by the change in money supply .

Best of Luck !! Keep Chegging !!